OpenStudy (anonymous):

A basketball team has a ratio of wins to losses of 3:1. After the team won six in a row, its ratio of wins to losses became 5:!. How many games had the team won BEFORE winning 6 games in a row?

OpenStudy (anonymous):

5:1 is the correction

OpenStudy (anonymous):

W/L = 3/1 (W+6)/L = 5/1 solve for W and L.